Steam is stored in a 0.3m3

rigid tank, initially at 7bar and 440oC. After a long 

time, steam temperature was measured and found to be 240oC. Find the final 

pressure as well as work and heat exchanged during that time.


1
Expert's answer
2021-07-06T11:37:20-0400

Gives


"T_1=440\u00b0C=713K,T_2=240\u00b0C=513K,P_1=7bar"

We know that


Pv=nRT

Case (1)

"P_1V_1=nRT_1"

Case (2)

"P_2V_2=nRT_2"

Gives

"V_1=V_2"

Both eqution we can written as

"\frac{P_1}{P_2}=\frac{T_1}{T_2}"

"P_2=\frac{T_2}{T_1}\times P_1"

"P_2=\frac{713}{513}\times7=9.72bar"